OpenStudy (anonymous):

A racing car has a uniform acceleration of 5 m/s². What distance will it cover in 20s after starting from rest?

OpenStudy (anonymous):

S=ut+1/2at2 S=? a=5m/s2 t=20s S=0+1/2∗5∗20∗20 S=1000m

OpenStudy (anonymous):

\[s=v _{0}t+1/2at ^{2}\] where s equals distance, v0 equals starting velocity, t equals elapsed time, and a equals acceleration. s= 0X20 + 0.5X5X20X20 = 0 + 1000 = 1000m or 1km

OpenStudy (anonymous):

\[s(t)=0,5 \times a \times t^{2} -> a=5 ; t=20 -> s=1 km\]